A mountain is 11,986 feet above sea​ level, and a valley is 263 feet below sea level. What is the difference in elevation between the mountain and the​ valley?
Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52781)

.

From the number of 11,986 subtract the number of -263.


If you will manage to subtract correctly, at the end you will get your answer.



Answer by math_tutor2020(3817)


Draw a vertical number line.
Mark point A at 11986
Point B is at 0 to represent sea level.
Point C is at -263
Don't worry about the diagram being to scale. A rough quick sketch will suffice.

The distance from A to B is 11986 feet.
The distance from B to C is 263 feet.

Add those values to get your final answer, which is the distance from A to C.

Why add when the textbook mentioned "difference"?
This is because what we've effectively done is A - C = 11986 - (-263) = 11986 + 263
The two negatives cancel out to form a positive.